The volcanic activity has been at a slightly lower level today, with rockfall signals once again being the predominant seismic triggers.
Unfortunately the broadband seismic system is still out of action today, and the short period system is still experiencing a back log of events for characterisation. Hence there are no earthquake counts for today. However, the number of triggers on the short period system today has been slightly lower than yesterday, and the predominant seismic activity has been long period earthquakes and rockfall signals.
Clear observations of the dome today showed that most of the rockfall activity was concentrated over the Galways area. There appeared to be no significant change in the northern face of the dome.
The GPS survey of the eastern network including Harris Lookout, Windy Hill, Hermitage, Long Ground and Lees Yard completed yesterday has now been processed. The line between Harris and Whites has shown 4 cm shortening between March and September this year followed by partial recovery. The line is now 2 cm shorter than its long term mean. The site at Hermitage continues its slow movement to the north-east (away from the volcano), and the total movement is now 10.7 cm of which 6.2 cm has occurred since September 20th 1997.
Air quality at all monitoring stations across the island is good today.
Although activity is mainly focused in the southern central part of the volcano, the area of growth on the dome can change very quickly. Growth activity appears now to be backing up behind the major build-up of material in Galways, and a renewal of rockfalls or pyroclastic flows down Gages or Tar River would not be a surprise in the near future.
Today, Dr. Willy Aspinall left the island after his month long tour of duty, and David "Patch" Silcott returned to Montserrat after completing his first semester at the University of the West Indies at St. Augustine, Trinidad. He will be helping out at the observatory over the Christmas holiday.